OpenStudy (anonymous):

Solve for x: (2/x-2) + (7/x^2-4) = (5/x) a. x = -(4/3) and x = 5 b. x = (4/3) and x = 5 c. x = (4/3 and x = –5 d. x = -(4/3) and x = –5 Thanks!

OpenStudy (bahrom7893):

2x(x+2)+7x=5(x-2)(2+2)

OpenStudy (bahrom7893):

2x^2+4x+7x=5x^2-20

OpenStudy (bahrom7893):

I'll let Lgba finish

OpenStudy (anonymous):

Okay thanks!!

OpenStudy (lgbasallote):

the lcm of everything here is (x+2)(x-2)x multiply everything by that...you'll get 2(xx+2) + 7x = 5(x^2-4) 2x^2 + 4x + 7x = 5x^2 -20 11x + 2x^2 = 5x^2 -20 3x^2 - 11x -20 = 0 (3x-4)(x+5) equate the factors to zero... 3x - 4 = 0; x +5 = 0 so 3x = 4 x = 4/3 and x = -5 darn bah was a fast typer @_@

OpenStudy (anonymous):

that was incorrect! i used that one & i got it wrong.